Boxes of​ Honey-Nut Oatmeal are produced to contain 15.0 ​ounces, with a standard deviation of 0.15 ounce. For a sample size of 49​, the ​3-sigma x (overbarx) chart control limits​ are:

Upper Control Limit ​(UCL overbar X) = __ ounces ​(round your response to two decimal​ places).

Lower Control Limit =(LCL=max
Answer #1

In the case of an Xbar chart, we have the following formula to find the UCL and LCL:

3 sigma UCL = Process mean + 3 * standard deviation / sqrt(n)
3 sigma LCL = Process mean - 3 * standard deviation / sqrt(n)

We have the following data:

Mean = 15
Standard deviation = 0.15
Sample size (n) = 49

UCL = 15 + (0.15 / SQRT(49) = 15 + 0.021 = 15.02
LCL = 15 - (0.15 / SQRT(49) = 15 - 0.021 = 14.98